Solve the differential equation by substitution suggested by equation. Show complete solution.


(dy/dx) = 2y/x + cos(y/x2)


1
Expert's answer
2021-10-06T15:26:11-0400

"\frac{dy}{dx}=\frac{2y}{x}+cos(\frac{y}{x^{2}})" this is equation (i)

Let y=Vx2

Then "\frac{dy}{dx}=2Vx+x^{2}\frac{dV}{dx}"

Substituting "\frac{dy}{dx}" in equation (i)

"2Vx+x^{2}\frac{dV}{dx}=2\frac{Vx^{2}}{x}+cos(\frac{Vx^{2}}{x^{2}})"

"x^{2}\frac{dV}{dx}=2Vx-2Vx+cos(V)"

"x^{2}\frac{dV}{dx}=cos(V)"

"\frac{dV}{cos(V)}=\frac{dx}{x^{2}}"

Integrating both sides;

"\int sec(V)dV=\int \frac{dx}{x^{2}}"

"ln(sec(V)+tan(V))=\frac{-1}{x}+C" this is equation (ii)

Substituting "V=\frac{y}{x^{2}}" into equation (ii)

"ln(sec(\frac{y}{x^{2}})+tan(\frac{y}{x^{2}}))=\frac{-1}{x}+C"